Session: Feb 2012 (session 1)
Wavelength: 6 cm
To give feedback on this experiment, click
(orfei at Wed Mar 07 13:00:29 2012) Success:
Hartebeesthoek (Jonathan Quick at Wed Mar 07 13:11:32 2012) Success:
No known problems
(Kaz Borkowski at Wed Mar 07 13:19:28 2012) Success:
No known problems.
(Geert Kuper at Wed Mar 07 13:46:45 2012) Failure: PHAS
Westerbork had a serious backend problem. A real time system and a control board of ADC unit 6 were broken. Only the last 2 minutes of this project
Yebes (Pablo de Vicente at Wed Mar 07 14:36:36 2012) Success:
No known problem
(Michael Lindqvist at Wed Mar 07 14:48:55 2012) Failure: WIND
To high wind, could not observe.
(Hua Zhang at Thu Mar 08 02:02:06 2012) Success:
(Bo Xia at Mon Mar 12 04:14:21 2012) Success:
(Uwe Bach at Mon Mar 12 11:26:15 2012) Success:
(Alastair Gunn at Mon Mar 19 15:20:16 2012) Minor failures:
Failed bias supply fixed by 3rd FTP scan.